OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2) A. (–2.5, 1.5) B. (–0.5, –0.5) C. (0.5, –0.5) D. (–3, –5)

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

mid point of two points say A(x1, y1) and B(x2, y2) is given as \[\frac{x1+x2}{2}, \ \ \ \ \ \ \ \ \ \ \frac{y1+y2}{2}\]

OpenStudy (ash2326):

Could you try now?

OpenStudy (anonymous):

no sorry dont understand

OpenStudy (ash2326):

midpoint is given by the average of the x coordinates and y coordinates here (x1, y1)= (-2, 1) and (x2, y2)=(-3, 2) \[\text{midpoint}= (\frac{-2+(-3)}{2}, \ \frac{1+2}{2})\]

OpenStudy (anonymous):

(D)

OpenStudy (anonymous):

am i right ????

OpenStudy (ash2326):

check again :)

OpenStudy (anonymous):

i dont know what to do

OpenStudy (ash2326):



OpenStudy (ash2326):

sum the x and y coordinates and divide by 2

OpenStudy (anonymous):

what is the formila

OpenStudy (anonymous):

sorry about the spelling

OpenStudy (ash2326):

\[\large \frac{x1+x2}{2}, \ \ \ \ \ \ \ \ \ \ \frac{y1+y2}{2}\]

OpenStudy (anonymous):

?? sorry but i dont know still =(

OpenStudy (anonymous):

(A) ??